OpenStudy (girlgoyleh.):

PLEASE HELP!! The value of y varies directly with x2, and y = 36 when x = 3. What is the value of y when x = 5?

OpenStudy (anonymous):

When you said "x2" do you mean "x squared"?

OpenStudy (girlgoyleh.):

yeah

OpenStudy (anonymous):

Refer to the attachment

OpenStudy (girlgoyleh.):

i still dont understand...

OpenStudy (anonymous):

Let y = k x^2 y=36 and x=3 Plug in the values for y and x into y = k x^2 and then solve the result for k. k = 4 y = 4 x^2 When x=5, y=4*5^2 y = 4*25 = 100

OpenStudy (girlgoyleh.):

thanks

OpenStudy (anonymous):

Your welcome and thank you for the medal.

OpenStudy (girlgoyleh.):

you're welcome :)